OpenStudy (anonymous):

SOLVE: y varies directly with x. When y = 20, x = 50. Find x when y = 36. Choose one answer. a. 75 b. 80 c. 90 d. 105

OpenStudy (anonymous):

y varies directly with x means: \[\large y = kx\] Now when y = 20, x = 50, by plugging in the values, we can find k here: \[20 = k \times 50 \implies k = \frac{2}{5}\]

OpenStudy (anonymous):

Now you have the value of k = 2/5 and y = 36, find x here using same equation that I have written above. Can you find?

OpenStudy (anonymous):

C...?